# Supervised heterogeneous network estimation via survival-based Bayesian graphical models

### v2

@@ 4 Data Analysis
We illustrate the proposed approach in constructing gene networks of different subgroups of cutaneous melanoma (SKCM) from The Cancer Genome Atlas (TCGA) data. Following existing literature [1], we consider the metastatic subjects and use the mRNA gene expression data and overall survival time for the predictors and response, respectively. $360$ subjects are available with $19,039$ gene expression measurements. Among them, $53.6\%$ are non-censored, with the observed survival times ranging from 2.6 to 357.1 months (median 44.05 months). As the sample size is relatively small, we follow the existing studies [12] and conduct gene filtering. Specifically, we first identify 2,260 genes with significance levels below $0.01$ from univariate Cox-regression models and then focus on those present in six pathways (pyrimidine metabolism, P53 signaling, cysteine and methionine metabolism, chemokine signaling, cell cycle, and hedgehog signaling pathways) that have demonstrated important implications for SKCM in the literature. With this preprocessing, 91 genes are obtained for downstream analysis.
@@
Figure 1: Data analysis: gene networks for the two subgroups identified by the proposed approach. In each network, the black edges are shared by the two subgroups, and the colored edges are unique in the corresponding subgroup.
Comparative analysis using the alternative approaches is also carried out with $K=2$. The summary results are reported in Table S7 (supplementary materials), which provide the number of subjects and network edges in the two subgroups identified by different approaches and their overlaps. The subgroups identified using different approaches are matched by correlation. In Table S7 (supplementary materials), different approaches are observed to identify subgroups with different subjects. To get a deeper understanding of the heterogeneity analysis performance, we examine whether these subgroups are associated with survival time. In Figure 2, we present the Kaplan–Meier (KM) curves of survival time of different subgroups identified by the proposed and alternative approaches. It is observed that the KM curves of the two subgroups are significantly distinguished from each other for the proposed approach (P value: 0.005). Comparatively, the FM approach shows less significant differences. BCGS and SCAN approaches fail to identify subgroups with distinct survival. These biologically reasonable findings indicate that the proposed approach can satisfactorily identify survival-related subgroups and also provide support for its effectiveness in clinical subgroup identification.
Figure 2: Data analysis: Kaplan-Meier curves for the two subgroups identified by different approaches.
In Table S7 (supplementary materials), different approaches demonstrate a moderate number of overlapping edges. To provide indirect support for the network estimation results, we use the resampling strategy and randomly divide the data into a training and a testing set. As in the literature [9], the values of the negative log-likelihood statistic (NLS) for the 91 genes are calculated to evaluate prediction accuracy. The average NLS values of 100 resamplings are 7102.118 for the proposed approach, 7658.492 for BCGS, 7674.739 for SCAN, 8044.482 for FMGemBag, and 8221.576 for FMJGL, which suggest a satisfactory network estimation accuracy of the proposed approach.
@@ 5 Discussion
In this study, we have employed the commonly used Gaussian mixture models for heterogeneity analysis, which show satisfactory performance. It can be of interest to explore alternative robust techniques, such as the exponential family graphical models proposed by [27] or other nonparametric graphical models, for studying heavy-tailed data. For survival modeling, we employ the Normal-error AFT model as a canonical example. Our framework is general and can be extended to more flexible survival models, such as the Weibull AFT model by appropriately modifying the likelihood function, representing a valuable direction for future research. Our focus in this data analysis has been mainly on the identification of gene networks using expression data. There is vast potential for a more comprehensive understanding of cancer mechanisms by incorporating other omics measurements, such as mutation and DNA methylation.
6 Disclosure statement
@@
7 Data Availability Statement
@@
Supplement to “Supervised Bayesian Joint Graphical Model for Simultaneous Network Estimation and Subgroup Identification”, including a graphical illustration of the method, algorithm details, sensitivity results, definition of $W$ in Assumption 2, technical proofs, simulation settings and results, and additional real data results.

### v4

@@ 5 Data analysis
We illustrate the proposed approach in constructing gene networks of different subgroups of cutaneous melanoma (SKCM) from TCGA data. Following existing literature (Ahmad and Fröhlich, 2017), we consider the metastatic subjects and use the mRNA gene expression data and overall survival time for the predictors and response, respectively. $360$ subjects are available with $19,039$ gene expression measurements. Among them, $53.6\%$ are non-censored, with the observed survival times ranging from 2.6 to 357.1 months (median 44.05 months). As the sample size is relatively small, we follow the existing studies (Hao et al., 2018) and conduct gene filtering. Specifically, we first identify 2,260 genes with significance levels below $0.01$ from univariate Cox-regression models and then focus on those present in six pathways (pyrimidine metabolism, P53 signaling, cysteine and methionine metabolism, chemokine signaling, cell cycle, and hedgehog signaling pathways) that have demonstrated important implications for SKCM in the literature. With this preprocessing, 91 genes are obtained for downstream analysis.
@@
Figure 3: Data analysis: gene networks for the two subgroups identified by the proposed approach. In each network, the highlighted edges are shared by the two subgroups.
Comparative analysis using the alternative approaches is also carried out with $K=2$. The summary results are reported in Table S4 (supplementary materials), which provide the number of subjects and network edges in the two subgroups identified by different approaches and their overlaps. The subgroups identified using different approaches are matched by correlation. In Table S4 (supplementary materials), different approaches are observed to identify subgroups with different subjects. To get a deeper understanding of the heterogeneity analysis performance, we examine whether these subgroups are associated with survival time. In Figure S13 (supplementary materials), we present the Kaplan-Meier (KM) curves of survival time of different subgroups identified by the proposed and alternative approaches. It is observed that the KM curves of the two subgroups are significantly distinguished from each other for the proposed approach (P value: 0.005). Comparatively, the FM approach shows less significant differences. BCGS and SCAN approaches fail to identify subgroups with distinct survival. These biologically reasonable findings indicate that the proposed approach can satisfactorily identify survival-related subgroups and also provide support for its effectiveness in clinical subgroup identification.
In Table S4 (supplementary materials), different approaches demonstrate a moderate number of overlapping edges. To provide indirect support for the network estimation results, we use the resampling strategy and randomly divide the data into a training and a testing set. As in the literature (Fan et al., 2019), the values of the negative log-likelihood statistic (NLS) for the 91 genes are calculated to evaluate prediction accuracy. The average NLS values of 100 resamplings are 7102.118 for the proposed approach, 7658.492 for BCGS, 7674.739 for SCAN, 8044.482 for FMGemBag, and 8221.576 for FMJGL, which suggest a satisfactory network estimation accuracy of the proposed approach.
@@ 6 Discussion
In this work, we employ a Bayesian hierarchical framework to accommodate heterogeneity at the molecular level and in molecular-clinical outcome associations. To enhance scalability, we compute MAP estimates via an EM algorithm rather than traditional MCMC. Theoretical guarantees are established based on these MAP estimates, whereas Bayesian uncertainty is quantified through PIPs, which further ensures network recovery consistency. This MAP plug‑in approach is well‑established in prior work, valued for its computational efficiency and solid theoretical foundation (Gan et al., 2019; Yang et al., 2021). Future directions include exploring MCMC-based approaches for potentially more accurate inference and improving computational efficiency in precision matrix estimation via node-wise strategies (Qin et al., 2024). We have employed the commonly used Gaussian mixture models for heterogeneity analysis. It can be of interest to explore alternative robust techniques, such as the exponential family graphical models or other nonparametric graphical models, for studying heavy-tailed data. In this study, we have used the Normal-error AFT model as a representative survival framework. Our proposed approach is general and could be extended to more flexible survival models—such as the Weibull AFT model—by adapting the likelihood accordingly. Furthermore, we have applied Laplace priors on $\beta_{k,j}$ and $\mu_{k,j}$ to maintain a balance between regularization and computational efficiency, since the primary focus here is network construction rather than variable selection among predictors. The spike-and-slab Laplace prior remains a promising alternative for future research that emphasizes predictor-level selection. Our focus in this data analysis has been mainly on the identification of gene networks using expression data. There is vast potential for a more comprehensive understanding of cancer mechanisms by incorporating other omics measurements, such as mutation and DNA methylation.
Acknowledgments
The authors thank the editors and reviewers for their invaluable feedback and insightful suggestions, which have significantly improved this paper.
Disclosure statement
@@ Acknowledgments
Data Availability Statement
@@
Supplement to “Supervised Bayesian Joint Graphical Model for Simultaneous Network Estimation and Subgroup Identification”, including a graphical illustration of the method, algorithm details, sensitivity results, assumptions, technical proofs, simulation settings and results, and additional real data results.